OpenStudy (anonymous):

Determine which binomial is a factor of -2x^3 + 14x^2 - 24x + 20 A.) x + 5 B.) x + 20 C.) x - 24 D.) x - 5

OpenStudy (anonymous):

its mcq so its easy to solve X is definitely wrong coz X is not the same as x

OpenStudy (anonymous):

if (x+5) is a factor, then x = -5 would make the equation 0

OpenStudy (anonymous):

if (x-5) is a factor, then x= 5 makes the equation 0

OpenStudy (anonymous):

substitute x = 5 and x = -5 see which one makes it 0

OpenStudy (anonymous):

Neither of them do.

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

There all 4 little x's I just didn't mean to make the two middle one's capital x's.

OpenStudy (anonymous):

So I'll try the other two to see which one is right

OpenStudy (anonymous):

i managed to find the answer by substituting a value

OpenStudy (anonymous):

try x = 5 again

OpenStudy (anonymous):

ok, I will

OpenStudy (anonymous):

When I put the equation like this -2(5)^3 + 14(5)^2 - 24(5) + 20 I get 100 + 1900 - 120 + 20 and when you add it up it equals 4900

OpenStudy (anonymous):

Is the answer x = 5 ?

hartnn (hartnn):

yes, x=5 is the root , so x-5 is the factor.

OpenStudy (anonymous):

So the answer is D ? x - 5

hartnn (hartnn):

yup.

OpenStudy (anonymous):

I wish I understood how you got that bc I tried to plugg 5 into where the x's are and I got the answer 4900 not 0

hartnn (hartnn):

5^3 = 125 125*2 = 250

hartnn (hartnn):

5^2 = 25 25*14=350

hartnn (hartnn):

24*5=120

OpenStudy (anonymous):

Ok, thanks :)

hartnn (hartnn):

got it ?